### v1

# Radiation properties and images of loop quantum Reissner-Nordström black hole with a thin accretion disk

Abstract
We investigate the characteristics of circular geodesics around loop quantum Reissner-Nordström black hole (LQRNBH) and the radiation properties and observational appearance of a thin accretion disk around it. By calculating the shadow radius and utilizing observational data from M87* and Sgr A*, we derive constraints on the quantum parameter $\zeta$ and charge parameter $Q$. The timelike circular geodesics around LQRNBH and the influence of the model parameters on the circular motion are also discussed. It is found that contrary to the case of the parameter $Q$, the increase of the quantum parameter $\zeta$ leads to the increase of the radius of the innermost stable circular orbit (ISCO). Then, by considering a thin accretion disk model, various radiation properties of the LQRNBH and the effects of the model parameters on them are studied. Concrete examples are provided for quantitative comparison of the radiation properties between LQRNBH and Schwarzschild black hole. With the ray-tracing method, the isoradial curves, redshift distributions, and the observed radiation fluxes of the direct and secondary images of the LQRNBH with the thin accretion disk for various model parameters and observation angles are numerically calculated and discussed. These results are beneficial for us to understand the physical consequences of loop quantum gravity effect.
I Introduction
Black holes (BHs), predicted by Einstein’s general relativity, are believed to be important astrophysical objects, which has successfully withstood the observation of gravitational waves since 2015 [1] and the observation of shadow images in recent years [2, 3]. However, it is known that black holes suffer from singularity problem which can not be avoided under certain common physical conditions in general relativity [4, 5], due to the geodesic incompleteness of spacetimes. Spacetime singularities may lead to the conflict between general relativistic dynamics and the unitarity of quantum evolution, i.e., BH information paradox [6, 7]. Thus, many efforts are made to resolve singularities of BHs, and typical approaches include string theory, loop quantum gravity, etc. (see [7] and references therein).
Symmetry-reduced cosmological models and BH models have been successfully constructed in loop quantum gravity framework [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], which resolve the classical singularities. A challenging issue in the effective theory of the quantum symmetry-reduced BH models is about the diffeomorphism covariance. Recently, this issue has been addressed through several approaches in spherically symmetric models, such as the quantum Oppenheimer-Snyder model [22, 23, 24] and loop quantum BHs (LQBHs) proposed in [25, 26] by constructing effective Hamiltonian constraints.
The LQBHs have been studied extensively from various aspects. The quasinormal modes of these spherical LQBHs have been explored by several recent works [27, 28, 29, 30, 31, 32, 33, 34]. The correspondence between grey-body factors and quasinormal modes of these LQBHs were also studied [35]. The features of the shadows of these quantum black holes were studied in [36, 37, 38, 39], and the constraints on the parameter encoding the quantum gravity effect from Event Horizon Telescope (EHT) observations [34, 40, 41] and gravitational wave observation [42] were also discussed. Furthermore, the optical appearance of the accretion disk around LQBHs and the strong gravitational lensing effects of the LQBHs were explored in [43, 44, 45, 41].
Recently, the framework developed in [25, 26] for constructing LQBHs has been extended to the electrovacuum setting with a cosmological constant [46]. By solving the effective dynamics determined by the effective Hamiltonians, the authors derived three LQBH models. When the cosmological constant is taken zero, solution I is just a Reissner-Nordström (RN) black hole with loop quantum gravity modification which will be abbreviated to LQRNBH. In this paper, we will investigate the optical appearance of the LQRNBH surrounded by a thin accretion disk, and explore the influence of the electric charge and quantum gravity parameter on the images of LQRNBH.
The remainder of this paper is organized as follows. In Section 2, we consider the shadow of the LQRNBH and constrain both the charge parameter and quantum parameter with the observational data of M87* and Sgr A*. In Section 3, we consider the timelike circular geodesics around the LQRNBH and discuss the influence of model parameters on the circular motion. In Section 4, we consider various radiation properties of the LQRNBH surrounded by a thin accretion disk and their dependence on the model parameters. In Section 5, we numerically study the isoradial curves, redshift factors and observed fluxes of the direct and secondary images of the LQRNBH with the thin accretion disk. The final section is devoted to the conclusion and discussion.
II The LQRNBH metric and Constraints on model parameters
The static and spherically symmetric LQRNBH with nonzero cosmological constant is proposed in [46]. In this work, we consider the asymptotically flat LQRNBH which can be obtained by setting $\Lambda=0$ in the general solution [46], whose line element is
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},$ (1)
$f(r)=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)\left(1+\frac{\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)\right),$ (2)
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the standard line element on a two-dimensional sphere in spherical coordinates, $M_{\mathrm{eff}}$ is defined as the effective mass, $Q$ is the charge, $\zeta$ is a quantum correction parameter which is proportional to the Planck length, $\gamma$ is the Barbero-Immirzi parameter, and $M$ is the mass of the black hole. In addition, we focus on the physically interesting case where $M\gg\zeta$, $Q<M$. This metric reduces to the solution obtained in [26, 25] when $Q\rightarrow 0$, and to the classical RN metric as $\zeta\rightarrow 0$.
Analysis of the function $f(r)$ shows that when $M\gg\zeta$ the LQRNBH has the same horizon structure as that for the corresponding RNBH. The radius of event horizon of the LQRNBH is the same as that of the corresponding RNBH and is unaffected by the quantum correction parameter $\zeta$, which implies that a distant observer can not distinguish LQRNBH from RNBH by the event horizon alone and more refined dynamical or radiative features are required. In the next, we consider the shadow radius of LQRNBH, which is affected by the quantum parameter $\zeta$, and the constraints on the LQRNBH parameters from the recent observation results of M87* and Sgr A*.
For the calculation of the shadow radius of a spherical black hole, many methods have been proposed and for a recent review we refer [47]. For a static and spherically symmetric black hole, the radius of the photon sphere $r_{ph}$ is determined by the largest positive root of the following equation [47]
$rf^{\prime}(r)-2f(r)=0.$ (3)
And, the shadow radius $R_{sh}$ of the black hole observed by an observer at infinity is
$R_{sh}=\frac{r_{ph}}{\sqrt{f(r_{ph})}}.$ (4)
For the LQRNBH, the equation satisfied by $r_{ph}$ is
$(r^{2}-3Mr+2Q^{2})(r^{4}+2Q^{2}\zeta^{2}-4Mr\zeta^{2}+2r^{2}\zeta^{2})=0.$ (5)
The root of the second factor of the above equation is in the interval $(M,2M)$, and the largest root of the first factor of the above equation is greater than $2M$. Thus, the radius of the photon sphere of LQRNBH is
$r_{ph}=\frac{3M}{2}+\frac{1}{2}\sqrt{9M^{2}-8Q^{2}}.$ (6)
Plugging this into (4), we can obtain the shadow radius $R_{sh}$ of the LQRNBH,
$R_{sh}=\frac{r_{ph}^{4}}{\sqrt{(r_{ph}^{2}-2Mr_{ph}+Q^{2})(r_{ph}^{4}+\zeta^{2}(r_{ph}^{2}-2Mr_{ph}+Q^{2}))}}.$ (7)
Although the event horizon and photon sphere is unaffected by the quantum parameter $\zeta$, the shadow radius depends on the parameter $\zeta$. When $Q\rightarrow 0$, the shadow radius reduces to $R_{sh}=\frac{27M^{2}}{\sqrt{27M^{2}+\zeta^{2}}}$, which is the same as the previous result in [41].
The Event Horizon Telescope (EHT) collaboration announced the images of two supermassive black holes: M87* at the center of the Messier 87 galaxy and Sgr A* at the center of the Milky Way. For M87*, the angular diameter of its shadow is $\theta_{\text{sh}}^{\text{M87*}}=(42\pm 3)\mu as$, its distance to Earth is $D_{\text{M87*}}=(16.8_{-0.7}^{+0.8})~\text{Mpc}$ and its estimated mass is $M_{\text{M87*}}=(6.5\pm 0.7)\times 10^{9}~M_{\odot}$ [2, 48, 49, 50]. With these data, the dimensionless shadow diameter of the black hole can be computed as $d_{\text{sh}}=\frac{D\theta_{\text{sh}}}{M}$, which is $d_{\text{sh}}^{\text{M87*}}=\frac{D_{\text{M87*}}~\theta_{\text{sh}}^{\text{M87*}}}{M_{\text{M87*}}}\simeq 11.0\pm 1.5.$ The corresponding shadow radius of the black hole, $R_{\text{sh}}^{\text{M87*}}$, is $4.75\lesssim R_{\text{sh}}^{\text{M87*}}\lesssim 6.25\,(1\sigma),$ which leads to following constraints on the model parameters of LQRNBH,
$0\leq\zeta\lesssim 2.304\,(1\sigma),~~0\leq Q\lesssim 0.681\,(1\sigma).$ (8)
For Sgr A*, the observed shadow angular diameter is $\theta_{\text{sh}}^{\text{Sgr A*}}=(48.7\pm 7.0)\mu as$, its distance to earth is $D_{\text{Sgr A*}}=(8.15\pm 0.15)~\text{kpc}$ and its mass is $M_{\text{Sgr A*}}=(4_{-0.6}^{+1.1})\times 10^{6}~M_{\odot}$ [3, 51, 52, 53, 54]. Besides these data, the EHT also provide the fractional deviation between the observed shadow radius and the shadow radius of a Schwarzschild BH, $\delta\equiv\frac{R_{sh}}{3\sqrt{3}M}-1$. Taking the average of the Keck- and VLTI-based estimates of $\delta$, the $1\sigma$ interval for $\delta$ is $0.125\lesssim\delta\lesssim 0.005~(1\sigma)$[55], which leads to the following constraints on the model parameters of LQRNBH,
$0\leq\zeta\lesssim 2.875\,(1\sigma),~~0\leq Q\lesssim 0.799\,(1\sigma).$ (9)
These two constraints on the parameters $\zeta$ and $Q$ are shown in Fig.(1). The red and blue curves are the boundaries of the constraints from M87* and Sgr A* respectively. When $Q=0$, the constraints on the quantum parameter $\zeta$ are consistent with the previous results [34, 41]. In the following analysis, we adopt the constraints from M87*.
Figure 1: Constraints on $\zeta$ and $Q$ from the observations of M87* and Sgr A*.
III Motion of massive test particles and Circular Orbits
In this section we consider the motion of a test neutral particle with mass $m$ in the LQRNBH spacetime and its circular orbits. The Lagrangian of the test particle is described as [56]:
$\mathcal{L}(x,\dot{x})=\frac{1}{2}mg_{\mu\nu}\dot{x^{\mu}}\dot{x^{\nu}}=\frac{1}{2}m\left(-f(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2})\right)$ (10)
where $\dot{x^{\mu}}=\frac{\mathrm{d}x^{\mu}}{\mathrm{d}\tau}$, and $\tau$ is the proper time (affine parameter) of the test particle. The canonical momentum of the particle can be written as
$p_{\mu}\equiv\frac{\partial\mathcal{L}}{\partial\dot{x^{\mu}}}=mg_{\mu\nu}\dot{x^{\nu}}.$ (11)
The explicit expressions of the four components of the canonical momentum of the particle are
$p_{t}=-mf(r)\dot{t},~p_{r}=m\frac{1}{f(r)}\dot{r},~p_{\theta}=mr^{2}\dot{\theta},~p_{\phi}=mr^{2}\sin^{2}\theta\dot{\phi}.$ (12)
The Killing vectors $\xi^{\mu}=(\frac{\partial}{\partial t})^{\mu}$ and $\eta^{\mu}=(\frac{\partial}{\partial\phi})^{\mu}$ are related with the conserved energy $E$ and the angular momentum $L$ of the test particle:
$E=-p_{t}\xi^{t}=mf(r)\dot{t},~~L=p_{\phi}\eta^{\phi}=mr^{2}\sin^{2}\theta\dot{\phi}.$ (13)
For a massive particle, its four velocity also satisfies the following normalization condition
$g^{\mu\nu}p_{\mu}p_{\nu}+m^{2}=0$ (14)
For timelike geodesics in the equatorial plane ($\theta=\frac{\pi}{2}$), using Eqs. (12)-(14), we can obtain the following first-order equations of motion:
$\dot{t}$ $=-\frac{1}{mf(r)}p_{t}=\frac{E}{mf(r)},$ (15)
$\dot{r}$ $=\frac{f(r)}{m}p_{r},$ (16)
$\dot{\phi}$ $=\frac{1}{mr^{2}\sin^{2}\theta}p_{\phi}=\frac{L}{mr^{2}\sin^{2}\theta}.$ (17)
According to the equivalence principle, the motion of the massive particle depends on the specific energy and the angular momentum, which are defined as $\tilde{E}=\frac{E}{m}$ and $\tilde{L}=\frac{L}{m}$, rather than their absolute values. So, in the following discussion of the paper, we will always use the specific energy and angular momentum unless otherwise stated. But for simplicity of symbols, we still use $E$ and $L$ to denote the specific energy and angular momentum.
From Eq. (16), we can obtain the radial equation of motion for the $r$-coordinate of the test particle
$\frac{\dot{r}^{2}}{2}+V_{\text{eff}}(r)=\frac{E^{2}}{2},$ (18)
where the effective potential $V_{\text{eff}}(r)$ is
$V_{\text{eff}}(r)=f(r)\left(1+\frac{L^{2}}{r^{2}}\right).$ (19)
The circular equatorial orbits of the test particle are determined by the conditions $V_{\text{eff}}(r)=\frac{1}{2}E^{2}$ and $\partial_{r}V_{\text{eff}}(r)=0$. From these conditions, we can obtain the specific energy $E$, the specific angular momentum $L$ and the angular velocity $\Omega$ of the particle on a circular equatorial orbit with radius $r$,
$\Omega$ $=\frac{\mathrm{d}\phi}{\mathrm{d}t}=\sqrt{-\frac{g_{tt,r}}{g_{\phi\phi,r}}}=\sqrt{\frac{f^{\prime}(r)}{2r}},$ (20)
$E$ $=-\frac{g_{tt}}{\sqrt{g_{tt}-g_{\phi\phi}\Omega^{2}}}=\sqrt{\frac{2f(r)^{2}}{2f(r)-rf^{\prime}(r)}},$ (21)
$L$ $=\frac{g_{\phi\phi}\Omega}{\sqrt{g_{tt}-g_{\phi\phi}\Omega^{2}}}=\sqrt{\frac{r^{3}f^{\prime}(r)}{2f(r)-rf^{\prime}(r)}}.$ (22)
In Fig. 2, we plot the energy $E$, the angular momentum $L$, and the angular velocity $\Omega$ of the particles on different circular orbits around LQRNBHs with various parameters. We can see that the distribution of the angular velocity $\Omega$ is subject to minimal influence from both parameters $Q$ and $\zeta$. For fixed value of quantum number $\zeta=1.0$, the increase of black hole charge $Q$ leads to the decrease of both the energy $E$ and angular momentum $L$, especially for particles on orbits near the ISCO. For fixed value of charge parameter $Q=0.3$, the energy $E$ increases while the angular momentum decreases as the increase of the quantum parameter $\zeta$. The influence of the quantum parameter $\zeta$ on $E$ is relatively obvious only for particles on orbits near the ISCO. The influence of the quantum parameter $\zeta$ on the angular momentum $L$ is much more obvious for particles in the far region. This can be understood by series expansions of $E$ and $L$ with respect to $r$. As $r\rightarrow\infty$, $E\sim 1-\frac{1}{2r}+\frac{3}{8r^{2}}+\frac{27-16Q^{2}}{16r^{3}}+\frac{675-576Q^{2}+64(Q^{4}+\zeta^{4})}{128r^{4}}+O(\frac{1}{r^{5}})$ and $L\sim\sqrt{r}-\frac{3-Q^{2}-\zeta^{2}}{2}\sqrt{\frac{1}{r}}+O(r^{-\frac{3}{2}})$. We can see that the leading order influences of $\zeta$ on $E$ and $L$ are $\frac{1}{r^{4}}$ and $\frac{1}{\sqrt{r}}$ respectively. Although $L$ is more sensitive to $\zeta$ in the region far from the event horizon, the influence of $\zeta$ on $L$ is combined with $Q$ in the leading order and next to leading order is needed to distinguish the influences from $Q$ and $\zeta$.
Figure 2: Energy $E$, angular momentum $L$, and angular velocity $\Omega$ of particles on circular orbits around a LQRNBH with various $\zeta$ (top row) and $Q$ (bottom row).
Figure 3: (a) Effective potential for particles on circular orbits as a function of the orbital radius. (b) Effective potential for a particle moving off the ISCO after small perturbation. Curves are plotted for various quantum parameter $\zeta$.
The stability of these circular orbits is determined by the second derivative of the effective potential. For a stable orbit, $\partial_{r}^{2}V_{\text{eff}}>0$. Usually, there exists an innermost stable circular orbit (ISCO) which is the boundary of the stable circular orbits. The ISCO is determined by an additional condition $\partial_{r}^{2}V_{\text{eff}}=0$, i.e.
$\partial_{r}^{2}V_{\text{eff}}=\frac{3L^{2}f(r)}{r^{4}}-\frac{2L^{2}f^{\prime}(r)}{r^{3}}+\frac{1}{2}(1+\frac{L^{2}}{r^{2}})f^{\prime\prime}(r)=0.$ (23)
In Fig.3(a), we show the effective potential for particles on circular orbit with radius $r$ for various values of the quantum parameter $\zeta$ and fixed charge parameter $Q=0.3$. The quantum parameter $\zeta$ increases the minimum value of the effective potential. In Fig.3(b), we show the potential function for particles getting off the ISCO. The potential energy undergoes a sharp decline inside the ISCO (i.e. the plunging region), leading particles to fall irreversibly into the black hole, and the quantum parameter $\zeta$ also increases the potential energy.
Simplifying Eq. (23) with Eqs.(21) and (22), we can obtain
$\begin{split}&r^{8}(-4Q^{4}+9Q^{2}r+(r-6)r^{2})+3r^{4}(Q^{2}+(-2+r)r)(-4Q^{4}+9Q^{2}r+(-6+r)r^{2})\zeta^{2}\\ &\quad-2(Q^{2}+(-2+r)r)^{2}[12Q^{4}+Q^{2}r(-33+8r)+r^{2}(24+r(-13+2r))]\zeta^{4}=0.\end{split}$ (24)
The largest real root of this equation is just the radius of the ISCO. In the limit $\zeta=0$ and $Q=0$, we can obtain $r=6$, i.e. $R_{I}=6M$, which is just the ISCO of the Schwarzschild BH. Numerically solving the above equation, the ISCO radius $R_{I}$ is shown as a function of $\zeta$ and $Q$ in Fig. 4. Taking into account the constraint (8), the minimal value of the ISCO radius of LQRNBH is $5.23M$. We can see that the radius of ISCO decreases smoothly as $Q$ increases, while an increase in $\zeta$ leads to a slight increase in the ISCO radius. This indicates that compared with the Schwarzschild BH case, the quantum gravity effect tends to push the ISCO slightly outward, while the charge of a black hole usually pulls the ISCO closer to the black hole.
Figure 4: The ISCO radius $R_{I}$ as a function of parameters $\zeta$ and $Q$.
IV Radiation Properties of the Thin Accretion Disk
In this section, we consider various radiation properties when the LQRNBH is surrounded by a geometrically thin accretion disk [57]. We will focus on the influence of the model parameters, especially the quantum parameter $\zeta$, on these properties.
In the process of a test particle falling from infinity to the black hole, the efficiency of converting its gravitational energy into radiation can be used as a key indicator to quantify the energy conversion capacity of the central black hole. Assuming all the radiation energy reaches to infinity and the falling particle at infinity is at rest ($E_{\infty}\approx 1$), the radiation efficiency $\epsilon$ of a black hole for a test particle falling from spatial infinity to the ISCO are defined as [58, 59]
$\epsilon=\frac{E_{\infty}-E_{\text{ISCO}}}{E_{\infty}}\approx 1-E_{\text{ISCO}}.$ (25)
In Table 1, we calculate the radiation efficiency of LQRNBH for various values of $Q$ and $\zeta$ under the constraint (8). It is obvious that the radius of the event horizon does not change with $\zeta$. It is found that the radiation efficiency decreases with the increasing $\zeta$ but increases when $Q$ increases.
Table 1: The event horizon radius $R_{h}$, ISCO radius $R_{I}$, ISCO energy $E_{I}$ and radiative efficiency $\epsilon$ for different values of $\zeta$ and $Q$.
$\zeta$ $Q$ $R_{h}$ $R_{I}$ $E_{I}$ $\epsilon$
1.5 0 2 6.05598 0.943554 5.645 %
1.5 0.1 1.99499 6.04110 0.943429 5.657 %
1.5 0.3 1.95394 5.92014 0.942387 5.761 %
1.5 0.5 1.86603 5.66670 0.940085 5.991 %
$\zeta$ $Q$ $R_{h}$ $R_{I}$ $E_{I}$ $\epsilon$
0 0.5 1.86603 5.60664 0.939166 6.083 %
0.5 0.5 1.86603 5.60747 0.939179 6.082 %
1 0.5 1.86603 5.61928 0.939359 6.064 %
1.5 0.5 1.86603 5.66670 0.940085 5.991 %
The thin accretion disk model considered in this work was proposed [57]. The radiative flux $\tilde{F}(r)$ of the disk is given by:
$\tilde{F}(r)=-\frac{\dot{M}}{4\pi\sqrt{-g/g_{\theta\theta}}}\frac{\Omega_{,r}}{(E-\Omega L)^{2}}\int_{R_{I}}^{r}(E-\Omega L)L_{,r}\,dr,$ (26)
where $\dot{M}$ is the mass accretion rate and $g$ is the determinant of the metric. For the convenience of comparison with astrophysical observation, we recover dimensions in the following discussion. The relation between the radiation flux $\tilde{F}$ and the dimensional radiation flux $F$ is [60, 61]
$F=\frac{c^{6}}{G^{2}M^{2}}\tilde{F}$ (27)
The values of physical parameters used here are: $c=2.99792458\times 10^{10}$ cm s-1, $M_{\odot}=1.989\times 10^{33}$ g, $G=6.67430\times 10^{-8}$ cm3 g-1 s-1.
In actual observation, due to the moving of gas in the disk and the strong gravitational effect of the black hole, the observed radiation flux by a far observer is affected by both Doppler and gravitational redshifts. The observed radiation flux can be expressed as [57, 62, 63, 64]
$F_{obs}=\frac{F}{(1+z)^{4}},$ (28)
where the redshift factor $1+z$ is defined as
$1+z=\frac{1+\Omega b\sin\theta_{0}\sin\alpha}{\sqrt{-g_{tt}-\Omega^{2}g_{\phi\phi}}},$ (29)
where $\theta_{0}$ is the inclination angle of the observer, $\alpha$ is the polar angle of the photon’s position on the observer’s imaging plane, and $b$ is the impact parameter of the photon relative to the line joining the source and the observer [62, 41]. Using the LQRNBH metric, the above redshift factor can be written explicitly as
$1+z=\frac{1+\frac{r^{4}(r-Q^{2})-\zeta^{2}(r^{2}-2r+Q^{2})(r^{2}-4r+3Q^{2})}{(r^{2}-2r+Q^{2})(r^{4}+\zeta^{2}(r^{2}-2r+Q^{2}))}\sin\theta_{0}\sin\alpha}{\sqrt{\frac{(2Q^{2}+(-3+r)r)(r^{4}+2Q^{2}\zeta^{2}-4r\zeta^{2}+2r^{2}\zeta^{2})}{r^{6}}}}.$ (30)
In Fig. 5, we show the radiation flux distribution of the accretion disk surrounding the black hole and the influence of parameters $Q$ and $\zeta$ on the flux. The mass accretion rate $\dot{M}$ is assumed to be $2\times 10^{-3}M_{\odot}/\text{yr}$ and the black hole mass is taken as $M=2\times 10^{6}M_{\odot}$. As $\zeta$ increases, the radiation flux decreases, which is consistent with the previous observation that the radiation efficiency decreases as $\zeta$ increases. In contrast, as $Q$ increases, both the peak value of the radiation flux and the corresponding radius $r_{m}$ of the peak increase.
Figure 5: The radiation flux $F$ of the thin accretion disk around the LQRNBH for various parameters $\zeta$ and $Q$.
When the accretion disk around the black hole is under local thermodynamic equilibrium, the local effective temperature $T_{\mathrm{eff}}$ of the accretion disk and the radiation flux satisfy the Stefan-Boltzmann relation $F(r)=\sigma T_{\mathrm{eff}}^{4}(r)$, the Stefan-Boltzmann constant $\sigma=\frac{2\pi^{5}k^{4}}{15c^{2}h^{3}}=5.67\times 10^{-5}$. Now we provide two examples to quantitatively compare the difference between LQRNBH and Schwarzschild black hole in several typical quantities.
For the first example, we adopt a mass of $M=4\times 10^{6}M_{\odot}$ and an accretion rate of $10^{-6}M_{\odot}/\text{yr}$, which refer the supermassive black hole at the center of our galaxy, Sgr A* [3, 51, 52, 53, 54]. In the case of a Schwarzschild black hole, the peak value of the radiative flux $F^{\max}$, peak value of the effective temperature $T_{\mathrm{eff}}^{\max}$, and the corresponding radial position $r_{m}$ are
$\begin{split}&F^{\text{max}}=2.21553\times 10^{12}\text{erg}~\text{cm}^{-2}\text{s}^{-1},\\ &T_{\mathrm{eff}}^{\text{max}}=1.40594\times 10^{4}K,\\ &r_{m}=9.53128[M].\end{split}$ (31)
The similar quantities for a typical LQRNBH with $\zeta=1.5$ and $Q=0.5$ are
$\begin{split}&F^{\max}=2.52334\times 10^{12}\text{erg}~\text{cm}^{-2}\text{s}^{-1},\\ &T_{\mathrm{eff}}^{\text{max}}=1.45242\times 10^{4}K,\\ &r_{m}=9.06422[M].\end{split}$ (32)
For the second example, we refer to observational data for the black hole M87* as our model parameters. We take the black hole mass $M=6.5\times 10^{9}M_{\odot}$, the accretion rate $\dot{M}=10^{-3}M_{\odot}/\mathrm{yr}$ [50, 65, 66]. The peak values of the radiative flux $F^{\max}$ and the local effective temperature $T_{\mathrm{eff}}^{\max}$, and the corresponding radial position $r_{m}$ of a Schwarzschild black hole are
$\begin{split}&F^{\max}=8.39016\times 10^{8}\text{erg}~\text{cm}^{-2}\text{s}^{-1},\\ &T_{\mathrm{eff}}^{\text{max}}=1.96128\times 10^{4}K,\\ &r_{m}=9.53128[M].\end{split}$ (33)
The similar quantities for a LQRNBH with $\zeta=1.5$ and $Q=0.5$ are
$\begin{split}&F^{\max}=9.55585\times 10^{8}\text{erg}~\text{cm}^{-2}\text{s}^{-1},\\ &T_{\mathrm{eff}}^{\text{max}}=2.02612\times 10^{4}K,\\ &r_{m}=9.06422[M].\end{split}$ (34)
(a) $\alpha=\frac{3\pi}{2}$
(b) $\alpha=\frac{\pi}{2}$
(c) $\alpha=\frac{3\pi}{2}$
(d) $\alpha=\frac{\pi}{2}$
Figure 6: The observed radiation flux $F_{\mathrm{obs}}$ in different directions ($\alpha=3\pi/2$ or $\pi/2$ ) on the observation plane for various parameters $\zeta$ and $Q$. The observation angle is $\theta_{0}=80^{\circ}$.
The observed flux $F_{\mathrm{obs}}$ of the LQRNBH accretion disk is shown in Fig. 6 for various values of $\zeta$ and $Q$. The mass accretion rate $\dot{M}$ is chosen as $2\times 10^{-3}M_{\odot}/\text{yr}$ and the black hole mass is taken as $M=2\times 10^{6}M_{\odot}$. The two subfigures in the left column (6(b), 6(d)) show the flux distribution in the most bright direction on the observation plane (i.e. $\alpha=\frac{3\pi}{2}$, particles in the disk approaching the observer), while the two subfigures in the right column (6(a), 6(c)) show the flux distribution in the most dark direction on the observation plane (i.e. $\alpha=\frac{\pi}{2}$, particles in the disk moving away from the observer). From Figs. 6(a) and 6(b), we see that the influence of $\zeta$ on the fluxes in the most bright and dark directions is different. The quantum parameter $\zeta$ lowers the observed flux in the bright direction while strengthening the observed flux in the dark direction. It is found in the previous discussion that the parameter $\zeta$ lowers the radiation flux of the accretion disk. However, the redshift factor is also affected by the parameter $\zeta$, which leads to the different influence of $\zeta$ on the observed fluxes in the bright and dark directions. From Figs. 6(c) and 6(d), we can see the charge parameter $Q$ has a consistent and obviously improving influence on the observed fluxes in both directions.
Figure 7: The maximum of the observed flux $F_{\text{obs}}^{\max}$ as functions of $\zeta$ and $Q$.
We also show the influence of the parameters $\zeta$ and charge $Q$ on the maximum of the observed radiation flux $F_{\text{obs}}^{\max}$ over the entire disk in Fig.7. The maximum of the observed radiation flux decreases monotonically as the increase of $\zeta$, while increasing monotonically as the increase of $Q$. This indicates the clear difference between the electromagnetic and quantum effects: electromagnetic effect enhances the radiation of the disk, but quantum effect decreases it.
V Optical Appearance and Image of the LQRNBH
In this section, we investigate the observational appearance of the LQRNBH surrounded by an optically thin accretion disk and discuss the influence of the parameters $\zeta$ and $Q$ on the images. Since the LQRNBH is spherically symmetric, the method used in this section is the backward ray-tracing approach which follows Refs. [62, 67, 41].
We first provide a brief review of the method. We assume the observer is far from the black hole and the observation plane is a flat plane with polar coordinates $(b,\alpha)$ which are introduced under Eq.(29). The coordinates of a point emitter on the accretion disk are $(r,\phi)$. The deflection angle of light traveling from its emitter to its image point on the observation plane is $\gamma$. Then, the trigonometric relation between $\alpha$ and $\gamma$ is
$\cot\gamma=\cos\alpha\tan\theta_{0},$ (35)
where $\alpha\in[0,2\pi]$, $\theta_{0}\in[0,\frac{\pi}{2}]$ and $\gamma\in[\frac{\pi}{2}-\theta_{0},\frac{\pi}{2}+\theta_{0}]$. On the trajectory plane of the light, its trajectory equation is
$(\frac{\mathrm{d}u}{\mathrm{d}\varphi})^{2}=\frac{1}{b^{2}}-u^{2}(1-2Mu+Q^{2}u^{2})(1+\zeta^{2}u^{2}(1-2Mu+Q^{2}u^{2})),$ (36)
where $u=\frac{1}{r}$, and $\varphi$ is the polar angle coordinate of a point on the trajectory. Here, $b$ is also the impact parameter of the light in the backward ray-tracing method. In order to obtain the image of a circular orbit in the disk (i.e. the isoradial curve in [62]), we need the relation between the coordinates $b$ and $\alpha$ of an image point in the isoradial curve on the observation plane, which can be numerically derived from the above trigonometric equation and trajectory equation with the deflection angle $\gamma$ serving as a bridge quantity.
Figure 8: Isoradial curves in the direct and secondary images of LQRNBH with a thin accretion disk at different observation angles. The curves from outside to inside correspond to $r=15M$, $r=10M$ and $r=6.2M$, respectively. Red, green and blue colors correspond to $\zeta=0,1,1.5$ or $Q=0,0.3,0.5$, respectively.
In Fig.8, we illustrate the direct and secondary images of three typical circular orbits on the accretion disk around the LQRNBH for various values of the parameters $\zeta$ and $Q$ and three different observation angles. In each subfigure, the radii of the three orbits from outside to inside are 15M,10M,6.2M15M,~10M,~6.2M, respectively. Images in the three columns from left to right correspond to three observation angles $17^{\circ}$, $53^{\circ}$, and $80^{\circ}$, respectively. The two figures in each column show the influences of the parameters $Q$ and $\zeta$ on the images. In each subfigure in the upper row, the quantum parameter $\zeta$ is fixed at $1.5$, and the charge parameter $Q$ takes the values 0, 0.3, 0.5, which are plotted in red, green, and blue curves, respectively. In each subfigure in the lower row, the charge parameter $Q$ is fixed at 0.5 and the quantum parameter $\zeta$ takes values 0, 1, 1.5, which are also plotted in red, green, and blue curves, respectively. As the parameter $\zeta$ or $Q$ increases, the isoradial curves always become smaller. The secondary images are more sensitive to the parameters than the direct images. It is also obvious that the larger the observation angle is, the more sensitive the images are to the parameters $\zeta$ and $Q$.
The gravitational redshift caused by the strong gravitational field of the black hole and the Doppler redshift caused by the relative motion between the particles in the accretion disk and the observer will have significant impact on the black hole images. Because the directions of the relative motion on both sides of the black hole are opposite, there are Doppler redshift and Doppler blueshift on either side of the black hole, respectively. When the Doppler blueshift exceeds the influence of gravitational redshift, there will be an overall blueshift for the image of some region in the disk.
In Fig.9, we show the influence of the charge parameter $Q$ on the redshift factor $z$ of direct and secondary images at different observation angles. The contours with $z=0,0.25,0.5,0.7,0.9$ are also plotted. The quantum parameter $\zeta$ is fixed at 1.5. The charge parameter $Q$ takes values 0 and 0.5. When the observation angle $\theta_{0}$ equals to zero, no Doppler redshift exists and the gravitational redshift of the disk is rotationally symmetric due to the spherical symmetry of the spacetime. As the increase of the observation angle, the Doppler redshift effect becomes remarkable, which results in the asymmetry of the images of the redshift factor. This can be seen from the three subfigures in each column in Fig.9. By comparing the first two subfigures in the first row, we can see that the innermost contour with $z=0.9$ moves outward with the increase of charge parameter, which indicates that the charge parameter makes the high-redshift region larger.
In Fig.10, we show the influence of quantum parameter $\zeta$ on the redshift factor $z$ of the direct and secondary images at different observation angles. Similarly, we also plot the redshift contours with $z=0,0.25,0.5,0.7,0.9$. By comparing the subfigures with different quantum parameters $\zeta$ in each row, we can see that the quantum parameter $\zeta$ reduces the value of redshift factor in the relatively high-redshift region (i.e. the righthand side of the black hole). This influence of the quantum parameter $\zeta$ is contrary to that of the charge parameter $Q$.
Figure 9: Influence of $Q$ on the direct and secondary images of the redshift factor $z$ of the accretion disk around the LQRNBH at different observation angles. Contours in the images from outside to inside correspond to $z=0,0.25,0.5,0.7,0.9$.
Figure 10: Influence of $\zeta$ on the direct and secondary images of the redshift factor $z$ of the accretion disk around the LQRNBH at different observation angles. Contours in the images from outside to inside correspond to $z=0,0.25,0.5,0.7,0.9$.
The direct and secondary images of observed radiation flux are shown in Fig.11 and Fig.12, which show the influence of the charge parameter $Q$ on the images and the influence of the quantum parameter $\zeta$ on the images, respectively. The mass of the LQRNBH is chosen as $M=2*10^{6}M_{\odot}$ and the accretion rate is chosen as $\dot{M}=2*10^{-3}M_{\odot}/\text{yr}$. The inner edge of the thin accretion disk is taken as the ISCO radius, $R_{I}$, and the outer edge is chosen as $15M$. In each figure, the observation angles of the three rows of subfigures from top to bottom are $\theta_{0}=80^{\circ}$, $53^{\circ}$ and $17^{\circ}$, respectively. Different flux contours are also shown to illustrate the influences of model parameters on the images. In the figures, the unit of the radiation flux is $10^{16}\text{erg}~\text{cm}^{-2}\text{s}^{-1}$. The values of the contour lines from outside to inside for the direct images at different observation angles are chosen as follows: the contour lines of the direct images at $\theta_{0}=80^{\circ}$ are $3.0,3.5,4.0,4.5,5.0,5.5$; the contour lines of the direct images at $\theta_{0}=53^{\circ}$ are $2.2,2.7,3.2,3.7,4.2$; the contour lines of the direct images at $\theta_{0}=17^{\circ}$ are $0.8,1.0,1.2,1.4$. The values of the contour lines from outside to inside for the secondary images at different observation angles are chosen as follows: for the secondary image at $\theta_{0}=80^{\circ}$, the values are $0.5,1,1.5,2,1.5$; for the secondary image at $\theta_{0}=53^{\circ}$, the values are $1,2$; for the secondary image at $\theta_{0}=17^{\circ}$, the values are $0.5,1$.
In Fig.11, the quantum parameter $\zeta$ is fixed at $1.5$. We show the influence of the charge parameter $Q$ on the observed radiation flux at different observation angles. From the two direct images at each observation angle, the contour coverage for the case with $Q=0.5$ is significantly larger than that with $Q=0$, and the maximum value of the observed radiation flux also increases, i.e. the charge parameter $Q$ increases the observed radiation flux remarkably. Similar conclusion applies to the secondary images. These can be understood from the facts that the charge parameter decreases the ISCO radius, i.e. increases the area of the accretion disk, and $Q$ also increases the radiative efficiency. We can also see that at small observation angle (e.g. $17^{\circ}$), the shape of the flux distribution of the direct image is almost a disk, and at large observation angle (e.g. $80^{\circ}$), the asymmetry of the shape of the flux distribution increases. In Fig.12, the charge parameter is fixed at $0.5$. We show the influence of the quantum parameter $\zeta$ on the observed radiation flux at different observation angles. By comparing the two direct images at each observation angle, we can see that the quantum parameter $\zeta$ reduces the maximum values of the observed flux and also reduces the coverage of the contour. The influence of the quantum parameter $\zeta$ on the direct images of the observed flux is more remarkable for the lager observation angle. Similar conclusion holds for the secondary images.
Figure 11: Influence of $Q$ on the direct and secondary images of observed radiation flux $F_{\text{obs}}$ at different observation angles. The unit of the flux is $10^{16}\text{erg}~\text{cm}^{-2}\text{s}^{-1}$.
Figure 12: Influence of $\zeta$ on the direct and secondary images of observed radiation flux $F_{\text{obs}}$ at different observation angles. The unit of the flux is $10^{16}\text{erg}~\text{cm}^{-2}\text{s}^{-1}$.
VI Conclusion and Discussion
In this study, we investigate the shadow, circular orbital dynamics, radiation properties and the images of the LQRNBH surrounded by a thin accretion disk, and systematically analyze the influence of model parameters $\zeta$ and $Q$ on the these quantities. We focus on the case where the black hole has the same horizon structure as the corresponding classical Reissner-Nordström black hole. The black hole shadow of LQRNBH is derived, and constraints on the quantum parameter $\zeta$ and charge $Q$ of the LQRNBH are analyzed with the observational data of M87* and Sgr A*, which provide the parameter space of our subsequent analyses.
The circular motion of massive particles around the LQRNBH is then considered and the effects of model parameters $\zeta$ and $Q$ on the energy $E$, angular momentum $L$, and angular velocity $\Omega$ are analyzed. It is found that the increase of charge parameter $Q$ leads to the decrease of the ISCO radius, but the increase of quantum parameter $\zeta$ leads to the increase of the ISCO radius. This is an interesting feature of the loop quantum gravity effect, which is different from the effect caused by the usual charge or rotation parameter. The influence of the model parameters on the distribution of the angular momentum of the particles on circular orbits remains obvious in the far region of the disk, providing a potential avenue for testing quantum gravity effect through measurement of the accretion disk.
The radiation efficiency $\epsilon$ increases as the increase of charge parameter $Q$, while it decreases slightly as the increase of quantum parameter $\zeta$, which indicates the competitive effects of electromagnetic and quantum gravity effects on the energy dissipation during the formation of the accretion disk around the LQRNBH. Then, we consider the influence of the parameters $\zeta$ and $Q$ on the local radiation flux of the accretion disk and the observed radiation flux. The local and observed radiation fluxes increase as the charge parameter $Q$ increases. The quantum parameter $\zeta$ reduces the local radiation flux, while its influence on the observed radiation flux varies with the direction $\alpha$ on the observation plane. For example, the quantum parameter $\zeta$ reduces the observed flux in the most bright direction ($\alpha=3\pi/2$) and increases the observed flux in the most bright direction ($\alpha=\pi/2$). Based on the observational data of M87* and Sgr A*, we also provide two concrete examples to quantitatively compare the difference between LQRNBH and Schwarzschild black hole in several typical quantities, such as the local radiation flux, temperature and location of their maximum values.
With the ray-tracing method, the direct and secondary images of the isoradial curves, redshift factors and observed fluxes are numerically studied. The effects of the model parameters and observation angles on these figures are illustrated. The direct and secondary images of the considered quantities are similarly affected by these factors. The Doppler redshift and shapes of the images are highly affected by the observation angles. At low observation angle (e.g. $17^{\circ}$), the redshift is dominated by the symmetric gravitational redshift and the shape of image of the disk is almost a disk. At high observation angle (e.g. $80^{\circ}$), the effect of Doppler redshift becomes remarkable, which leads to the asymmetry of the redshift distribution around the black hole, and the image of the disk is shaped like a straw hat. An increase in $Q$ leads to the increase of the maximum value of the observed flux and enlarges the coverage of contours in the flux image. The quantum parameter $\zeta$ makes the distribution of the observed flux more uniform, i.e. it reduces the maximum value of the observed flux and increases the minimum value of the observed flux. The larger the observation angle, the more obvious the effects of $\zeta$ and $Q$ on the direct and secondary images of the redshift factors and observed fluxes.
The results of LQRNBH in this work provide several observational signals that can be used to distinguish it from the classical RNBH. However, our analysis focuses on the asymptotically flat LQRNBH and thin disk model. Subsequent research can be extended to the LQRNBH with nonzero cosmological constants and LQRNBH surrounded by other disk models [68, 69], or dynamical accretion disk [70]. It is also interesting to explore the features of the quasinormal modes of the LQRNBH. These further research may yield more insights on our understanding of the quantum gravity effects.